# TabImpute: Accurate and Fast Zero-Shot Missing-Data Imputation with a Pre-Trained Transformer

### v1

@@ Abstract
Missing data is a pervasive problem in tabular settings. Existing solutions range from simple averaging to complex generative adversarial networks. However, due to huge variance in performance across real-world domains and time-consuming hyperparameter tuning, no default imputation method exists. Building on TabPFN, a recent tabular foundation model for supervised learning, we propose TabImpute, a pre-trained transformer that delivers accurate and fast zero-shot imputations requiring no fitting or hyperparameter tuning at inference-time. To train and evaluate TabImpute, we introduce (i) an entry-wise featurization for tabular settings, which enables a $100\times$ speedup over the previous TabPFN imputation method, (ii) a synthetic training data generation pipeline incorporating realistic missingness patterns, which boosts test-time performance, and (iii) MissBench, a comprehensive benchmark for evaluation of imputation methods with $42$ OpenML datasets and $13$ missingness patterns. MissBench spans domains such as medicine, finance, and engineering, showcasing TabImpute’s robust performance compared to $11$ established imputation methods.
@@ 1 Introduction
Missing data is ubiquitous across tabular datasets, affecting statisticians, economists, health officials, and businesses. For example, healthcare datasets may lack some recorded blood pressure measurements, or datasets merged from multiple sources may only share partial features. Regardless of the source, missing data must be imputed to numerical values before employing statistical or machine learning models.
@@
Figure 1: Evaluation on real-world OpenML data: MissBench. We compare TabImpute and TabImpute+ (ensembled method) with $11$ other popular methods on MissBench. In panel (a), we plot the imputation accuracy (defined as 1 - normalized RMSE), which is calculated for each method, normalized within a dataset, and averaged across datasets and $13$ missingness patterns. Error bars indicate 95% confidence intervals. In panel (b), we compare the runtime per table entry. Any method not labeled (GPU) is tested on a CPU because that method is not GPU-compatible. TabPFN on CPU is significantly slower, so we do not include it. See Sec. 3 for our exact computing specifications and Sec. 4 for accuracy score methodology.
@@
Rubin (1976) proposed analyzing missingness based on its relationship with the data-generating process to determine whether missingness biases downstream estimation. Rubin demonstrated that when missingness operates independently of the underlying data, the observed data distribution provides an unbiased foundation for estimation. This framework categorizes missingness into three classes: Missing Completely At Random (MCAR), Missing At Random (MAR), and Missing Not At Random (MNAR) (Van Buuren, 2012; Sportisse et al., 2020a; b). MCAR defines scenarios where missingness is independent of observed and missing values; MAR denotes settings where missingness depends on observed variables; and MNAR—most general of them all—denotes settings where missingness can arbitrarily depend on observed and unobserved variables.
Imputation methods range from simple approaches (e.g., constant default values and averages) to more sophisticated techniques like nearest-neighbor-based methods (Batista & Monard, 2003) and random forest regression, notably the MissForest algorithm (Stekhoven & Bühlmann, 2011). However, each method in the literature is typically tailored for specific settings, with performance varying significantly across datasets, domains, and missingness patterns (Van Buuren, 2012; Jarrett et al., 2022; Agarwal et al., 2023; Ibrahim et al., 2005). Another class of methods come from matrix completion including USVT (Chatterjee, 2015), SoftImpute (Hastie et al., 2015), and nearest neighbor-based methods (Agarwal et al., 2023; Dwivedi et al., 2022; Chin et al., 2025). While USVT and SoftImpute were typically designed for MCAR settings, nearest neighbor-based methods are designed for complex MNAR settings (Chin et al., 2025).
@@
Building on recent advances in tabular representation learning (Hollmann et al., 2023; Ye et al., 2025), we propose a pre-trained transformer specifically designed for the tabular missing-data problem that produces accurate and fast zero-shot imputations. TabPFN is a pre-trained transformer model for supervised learning that performs well across a variety of domains without any fine-tuning (Hollmann et al., 2025). The team behind TabPFN created an imputation method in their tabpfn-extensions Python package by using the TabPFN model in iterative column-wise imputation (available on GitHub [^1] https://github.com/PriorLabs/tabpfn-extensions/blob/main/src/tabpfn_extensions/unsupervised/unsupervised.py).
Starting point of this work is a comprehensive evaluation of the pre-existing methods, and we do so by building a new benchmark MissBench–consisting several real-world OpenML datasets and realistic missingness patterns induced on them. Next, we leverage the TabPFN architecture to introduce a new method TabImpute, which achieves state-of-the-art-performance on this benchmark. More specifically, our contributions can be summarized as follows:
We improve on this approach by introducing a new entry-wise featurization (EWF), allowing parallel prediction of each missing value using TabPFN’s model, denoted EWF-TabPFN. To further improve on TabPFN’s model for missing data imputation, we train a new underlying model, TabImpute, to better fit this class of tasks, outperforming fine-tuning TabPFN (shown in Sec. 4). Finally, we adaptively ensemble EWF-TabPFN’s and TabImpute’s predictions to achieve state-of-the-art results in imputation accuracy on our comprehensive benchmark. We denote the final ensembled method as TabImpute+.
@@
The main contributions of this work can be summarized as follows:
  • We introduce a novel, comprehensive test bench, denoted MissBench, using $42$ real-world OpenML datasets and $13$ missingness patterns (see Sec. 4 for details). Our benchmark builds on previous work, namely HyperImpute (Jarrett et al., 2022) comprising of $13$ UC Irvine (UCI) datasets (Kelly et al., 2024), a subset of OpenML, with $3$ missingness patterns, and GAIN (Yoon et al., 2018) comprising of $6$ UCI datasets with $1$ missingness pattern.
  • Next we introduce a state-of-the-art method in following steps:
    • First, we propose a new entry-wise missing data featurization (see Sec. 3.1 for details). Notably, this new featurization, when directly used with the pre-trained TabPFN model of Hollmann et al. (2023), provides a significant improvement in speed and accuracy over TabPFN. We call this method EWF-TabPFN.
    • Next, we develop a synthetic data generation pipeline to create training datasets with missing values covering a wide range of missingness patterns (see Sec. 3.2, Sec. A.2, and Tab. 6 for details). We pre-train a TabPFN architecture on this pipeline with our new featurization—we call this method TabImpute. This model provides an improvement over EWF-TabPFN across several missingness patterns.
    • Finally, we introduce an adaptive ensembling imputation method, denoted TabImpute+, to combine the power of TabImpute and EWF-TabPFN (see Sec. 3.4 for details). This allows us to leverage the underlying world-knowledge of EWF-TabPFN and domain-knowledge for missing data from TabImpute to yield accurate imputations across heterogeneous datasets.
Our code and implementation details for all our contributions above can be accessed on GitHub. [^2] https://github.com/jacobf18/tabular
Figure 3: Overview of our contributions. The first row demonstrates TabPFN’s imputation method, which performs iterative column-by-column imputation. We build on this by introducing an entry-wise featurization, as shown in the second row. We create a new synthetic data-generator for missingness data to train our model, TabImpute, shown in green (Sec. 3.2 and Sec. 3.3, respectively). Lastly, we ensemble TabImpute with TabPFN’s model using our features to create TabImpute+ (Sec. 3.4). We adaptively evaluate all the imputers on the comprehensive and rich set of OpenML datasets with many missingness patterns applied (Sec. 4).
@@ 1.1 Previous Work
We build primarily on missing data imputation and tabular representation learning (TRL). While missing data imputation is well-studied with established theory dating back to the 1970s (Rubin, 1976), TRL using tabular foundation models is relatively new (Müller et al., 2022; Zhang et al., 2025; Hollmann et al., 2023; 2025). Below, we describe relevant work that we directly compare against or build upon.
@@
Given the widespread nature of missing data, numerous imputation techniques have been proposed. These include averages, linear models over columns (Efron, 1994), random forest models (Hindy et al., 2024), ensemble methods (Jarrett et al., 2022), nearest neighbor-based methods (Chin et al., 2025), and even using generative adversarial networks (Yoon et al., 2018). For fully numerical data, matrix completion methods like SoftImpute (Hastie et al., 2015) have also been employed.
HyperImpute (Jarrett et al., 2022) combines the power of multiple classical imputation methods through iterative imputation (IM). IM loops over each column, using other columns to predict missing values until convergence. HyperImpute optimizes the imputer at each iteration over candidate methods, including MICE (Royston & White, 2011), SoftImpute (Hastie et al., 2015), column mean, MissForest (Stekhoven & Bühlmann, 2011), and Optimal Transport (Muzellec et al., 2020). While achieving state-of-the-art performance across diverse missingness patterns and supporting categorical variables, HyperImpute is designed for supervised learning settings and cannot handle matrices with entirely missing columns (common in causal inference and panel data (Agarwal et al., 2023)). Additionally, optimization requires significantly more time than base methods.
@@
Tabular representation learning focuses on building models that can generalize across diverse tabular domains. The pioneering work in this area is TabPFN, a tabular foundation model for supervised learning. TabPFN was proposed in Hollmann et al. (2023) and subsequently improved in Hollmann et al. (2025). Since TabPFN’s introduction, numerous variants have emerged to address scalability and performance limitations, as well as further work aimed at clarifying its internal representations (Zhang et al., 2025; Ye et al., 2025). Recent advances include TabICL (Qu et al., 2025), a scalable foundation model that extends supervised learning capabilities to datasets with up to 500K samples through a novel two-stage architecture with column-then-row attention mechanisms. Other notable models include MITRA (Zhang & Robinson, 2025), a tabular foundation model pre-trained on purely synthetic data from a mix of random classifiers/regressors, and Toto (Cohen et al., 2025), which is optimized for time series forecasting on observability metrics. Additionally, CausalFM (Ma et al., 2025) enables Bayesian causal inference through structural causal model priors similar to TabPFN, while DO-PFN (Robertson et al., 2025) extends the PFN framework to estimate counterfactual distributions from observational data.
@@ 2 Background on Prior-data Fitted Networks and TabPFN
Prior-data Fitted Networks (PFNs) are a class of models that learn to approximate Bayesian inference for a given prior (Müller et al., 2022). Instead of fitting a new model from scratch, a PFN is an individual, large pre-trained Transformer to perform classification or regression in a single forward pass. This process, known as in-context learning (ICL), allows the model to make predictions using sequences of labeled examples provided directly in the input, without requiring any gradient updates (Dong et al., 2024). The entire prediction algorithm is contained in the weights of the network, which is trained once on millions of synthetically generated datasets sampled from the prior. At inference time, the trained PFN takes a real-world dataset, composed of training and test samples, as a set-valued input and returns a distribution over the output space. This output space is categorical for classification tasks and the real line for regression tasks.
@@
PFNs are rooted in Bayesian supervised learning, where the primary objective is to model the posterior predictive distribution (PPD) (MacKay, 1992; Seeger, 2004; Blei et al., 2017). Since computing the PPD is often intractable (MacKay, 1992), PFNs instead learn to approximate the PPD offline through a process called synthetic prior fitting Müller et al. (2022). This is achieved using a prior specified by a sampling scheme that first samples a data-generating mechanism, $\phi\sim p(\phi)$, and then samples a synthetic dataset, $D\sim p(D|\phi)$. This process is repeated to generate millions of diverse datasets for training. The network’s parameters, $\theta$, are then optimized to predict held-out test samples ($D_{\text{test}}\subset D$) conditioned on the rest of the dataset ($D_{\text{train}}=D\setminus D_{\text{test}}$). The training objective is to minimize the negative log likelihood (NLL) loss on these held-out examples, averaged over the prior distribution of datasets $p(D)$:
@@
TabPFN is a PFN built specifically for tabular supervised learning (Hollmann et al., 2023). The model employs a novel two-way attention mechanism specifically designed for tabular data. Unlike standard transformers (Vaswani et al., 2023) that treat tabular data as sequential tokens, TabPFN assigns separate representations to each cell in the table. The architecture uses alternating attention patterns: each cell first attends to other features within its row (inter-feature attention), then attends to the same feature across all rows (inter-sample attention). This design ensures permutation invariance for both samples and features while enabling efficient scaling to larger tables than those seen during training. TabPFN v2 (Hollmann et al., 2025) retains the core training paradigm of the original TabPFN while introducing several key enhancements that improve accuracy, runtime, and applicability. Going forward, when referring to TabPFN, we mean TabPFN v2 since it is the most up-to-date version.
@@ 3 Training TabImpute on Synthetic Data
We develop a featurization for tabular missing data imputation that enables better utilization of TabPFN’s architecture, create a synthetic data generation pipeline across diverse missingness patterns, and employ an adaptive training algorithm to balance performance across all patterns. Training used $8$ H200 GPUs and an Intel Xeon Platinum 8592+ CPU over approximately one week, processing $25$ million synthetic tables. Our model matches TabPFN’s size and runs on CPU-only systems. Evaluation used $1$ H200 GPU.
@@ 3.1 Entry-wise Featurization and Architecture
We use TabPFN’s base architecture with one modification: removing the attention mask to allow training points to attend to test points. Since our train/test set sizes vary randomly with missingness patterns (unlike TabPFN’s controlled synthetic generation), we remove the mask to enable parallel batch training. TabPFN’s mask prevents train points from seeing test feature distributions, which is important in general supervised learning. However, our test set is created using data already available to the observed points, thus alleviating any data-leakage concerns.

### v3

@@ Abstract
Missing data is a pervasive problem in tabular settings. Existing solutions range from simple averaging to complex generative adversarial networks, but due to each method’s large variance in performance across real-world domains and time-consuming hyperparameter tuning, no default imputation method exists. Building on TabPFN, a recent tabular foundation model for supervised learning, we propose TabImpute, a pre-trained transformer that delivers accurate and fast zero-shot imputations requiring no fitting or hyperparameter tuning at inference-time. To train and evaluate TabImpute, we introduce (i) an entry-wise featurization for tabular settings, which enables a $100\times$ speedup over the previous TabPFN imputation method, (ii) a synthetic training data generation pipeline incorporating realistic missingness patterns, and (iii) MissBench, a comprehensive benchmark with $42$ OpenML datasets and $13$ new missingness patterns. MissBench spans domains such as medicine, finance, and engineering, showcasing TabImpute’s robust performance compared to $12$ established imputation methods.
@@ 1 Introduction
Missing data is ubiquitous across tabular datasets, affecting statisticians, economists, health officials, and businesses. For example, healthcare datasets may lack some recorded blood pressure measurements, or datasets merged from multiple sources may only share partial features. Regardless of the source, missing data must be imputed to numerical values before employing statistical or machine learning models. Imputation methods range from simple approaches (e.g., constant values and averages) to more sophisticated techniques like nearest-neighbor-based methods (Batista01052003), matrix factorization approaches such as SoftImpute (hastie2015matrix), and random forest regression, notably the MissForest algorithm (missforest2011). However, each method in the literature is typically tailored for specific settings, with performance varying significantly across datasets, domains, and missingness patterns (van2012flexible; jarrett2022hyperimpute; agarwal2023causal; Ibrahim01032005). Building on recent advances in tabular representation learning (hollmann2023tabpfn; ye2025closer), we propose a pre-trained transformer specifically designed for the tabular missing-data problem that produces accurate and fast zero-shot imputations.
@@
Figure 1: Evaluation on real-world OpenML data: MissBench. We compare TabImpute with $12$ other popular methods on MissBench. In panel (a), we plot the imputation accuracy (defined as 1 - normalized RMSE), which is calculated for each method, normalized within a dataset, and averaged across datasets and $13$ missingness patterns. Error bars indicate 95% confidence intervals. In panel (b), we compare the runtime per table entry. Any method not labeled (GPU) is tested on a CPU because that method is not implemented for GPUs. See Sec. 3 for our exact computing specifications and Sec. 4 for accuracy score methodology.
@@
rubin1976inference proposed analyzing missingness based on its relationship with the data-generating process to determine whether missingness biases downstream estimation. Rubin demonstrated that when missingness operates independently of the underlying data, the observed data distribution provides an unbiased foundation for estimation. This framework categorizes missingness into three classes: Missing Completely At Random (MCAR), Missing At Random (MAR), and Missing Not At Random (MNAR) (van2012flexible; sportisse2020aimputation; sportisse2020b). MCAR defines scenarios where missingness occurs uniformly and independently of all data values. MAR encompasses cases where missingness depends on observed variables that can be appropriately conditioned upon during analysis. MNAR describes situations where missingness depends on unobserved factors that cannot be easily conditioned on.
While this framework characterizes how missingness relates to the underlying data-generating process, it largely treats the missingness indicators for each variable as conditionally independent across entries. Recent work on structured missingness argues that, especially in large-scale multi-source datasets, missingness itself can exhibit rich multivariate structure that is not captured by the standard MCAR/MAR/MNAR taxonomy. In particular, mitra2023learning introduce structured missingness as an umbrella term for mechanisms in which missingness follows systematic patterns. jackson2023completecharacterisationstructuredmissingness provide a complementary characterization that allows dependencies among missingness indicators for different variables.
Previous work typically introduced new missingness patterns within these categories and proposed pattern-specific solutions. For instance, agarwal2023causal proposed a nearest neighbor-based matrix completion method specifically designed for a block-wise MNAR pattern. We instead develop a single method that performs well across diverse patterns and data domains by building on TabPFN, a popular tabular foundation model for supervised learning.
@@
TabPFN is a pre-trained transformer model for supervised learning that performs well across a variety of domains without any fine-tuning (hollmann2025accurate). The team behind TabPFN created an imputation method in their tabpfn-extensions Python package by using the TabPFN model in iterative column-wise imputation (available on GitHub [^1] https://github.com/PriorLabs/tabpfn-extensions/blob/main/src/tabpfn_extensions/unsupervised/unsupervised.py). However, when evaluated on our new benchmark MissBench–consisting of $42$ real-world OpenML datasets and $13$ missingness patterns–TabPFN’s imputation approach shows low accuracy and slow runtime. We improve on this approach by introducing a new entry-wise featurization (EWF), allowing parallel prediction of each missing value using TabPFN’s model, denoted EWF-TabPFN. To further improve on EWF-TabPFN, we train a new underlying model specifically designed for tabular data imputation, TabImpute, to better fit this class of tasks, achieving the imputation accuracy shown in Fig. 1 and Tab. 1.
@@
The main contributions of this work can be summarized as follows (also shown in Fig. 3):
  • We propose a new state-of-the-art pre-trained transformer model, TabImpute, for missing data imputation by building on TabPFN and introducing a new entry-wise missing data featurization (see Sec. 3.1 for details).
  • We develop a synthetic data generation pipeline to create training datasets with a comprehensive collection of missing values covering a wide range of MNAR patterns (see Sec. 3.2, Sec. A.2, and Tab. 10 for details).
  • We demonstrate TabImpute’s performance on a novel, comprehensive test bench, denoted MissBench, using $42$ real-world OpenML datasets and $13$ missingness patterns (see Sec. 4 for details). Our benchmark builds on previous work by testing more datasets and more missingness patterns: HyperImpute (jarrett2022hyperimpute) tests on $13$ UC Irvine (UCI) datasets (Kelly2024UCI), a subset of OpenML, with $3$ missingness patterns, and GAIN (gain2018) tests on $6$ UCI datasets with $1$ missingness pattern.
Our code and implementation details for all our contributions above can be accessed on GitHub. [^2] https://anonymous.4open.science/r/tabular-6F65/README.md
Figure 3: Overview of our contributions. The first row demonstrates TabPFN’s imputation method, which performs iterative column-by-column imputation. We build on this by introducing an entry-wise featurization, as shown in the second row. We create a new synthetic data-generator for missingness data to train our model, TabImpute, shown in green (Sec. 3.2 and Sec. 3.3, respectively). We evaluate all the imputers on the comprehensive and rich set of OpenML datasets with many missingness patterns applied (Sec. 4).
@@ 1.1 Previous Work
We build primarily on missing data imputation and tabular representation learning (TRL). While missing data imputation is well-studied with established theory dating back to the 1970s (rubin1976inference), TRL using tabular foundation models is relatively new (muller2022transformers; zhang2025tabpfnmodelruleall; hollmann2023tabpfn; hollmann2025accurate). Below, we describe relevant work that we directly compare against or build upon.
@@
Given the widespread nature of missing data, numerous imputation techniques have been proposed. These include averages, linear models over columns (efron1994missing), random forest models (missforest_python), ensemble methods (jarrett2022hyperimpute), nearest neighbor-based methods (chin2025n), and even using generative adversarial networks (gain2018). For fully numerical data, matrix completion methods like SoftImpute (hastie2015matrix) have also been employed.
HyperImpute (jarrett2022hyperimpute) combines the power of multiple classical imputation methods through iterative imputation (IM). IM loops over each column, using other columns to predict missing values until convergence. HyperImpute optimizes the imputer at each iteration over candidate methods, including MICE (royston2011multiple), SoftImpute (hastie2015matrix), column mean, MissForest (missforest2011), and Optimal Transport (muzellec2020missing). While achieving great imputation accuracy across diverse missingness patterns and supporting categorical variables, HyperImpute is designed for supervised learning settings and cannot handle matrices with entirely missing columns (common in causal inference and panel data (agarwal2023causal)). Additionally, optimization requires significantly more time than base methods.
@@
Tabular representation learning focuses on building models that can generalize across diverse tabular domains. The pioneering work in this area is TabPFN, a tabular foundation model for supervised learning. TabPFN was proposed in hollmann2023tabpfn and subsequently improved in hollmann2025accurate. Since TabPFN’s introduction, numerous variants have emerged to address scalability and performance limitations, as well as further work aimed at clarifying its internal representations (zhang2025tabpfnmodelruleall; ye2025closer). Recent advances include TabICL (qu2025tabicl), a scalable foundation model that extends supervised learning capabilities to datasets with up to 500K samples through a novel two-stage architecture with column-then-row attention mechanisms. Other notable models include MITRA (MITRA), a tabular foundation model pre-trained on purely synthetic data from a mix of random classifiers/regressors, and Toto (cohen2025timedifferentobservabilityperspective), which is optimized for time series forecasting on observability metrics. Additionally, CausalFM (ma2025foundationmodelscausalinference) enables Bayesian causal inference through structural causal model priors similar to TabPFN, while DO-PFN (robertson2025dopfnincontextlearningcausal) extends the PFN framework to estimate counterfactual distributions from observational data.
@@ 2 Background on Prior-data Fitted Networks and TabPFN
Prior-data Fitted Networks (PFNs) are a class of models that learn to approximate Bayesian inference for a given prior (muller2022transformers). Instead of fitting a new model from scratch, a PFN is an individual, large pre-trained Transformer to perform classification or regression in a single forward pass. This process, known as in-context learning (ICL), allows the model to make predictions using sequences of labeled examples provided directly in the input, without requiring any gradient updates (dong2024surveyincontextlearning). The entire prediction algorithm is contained in the weights of the network, which is trained once on millions of synthetically generated datasets sampled from the prior. At inference time, the trained PFN takes a real-world dataset, composed of training and test samples, as a set-valued input and returns a distribution over the output space. This output space is categorical for classification tasks and the real line for regression tasks.
@@
PFNs are rooted in Bayesian supervised learning, where the primary objective is to model the posterior predictive distribution (PPD) (mackay1992practical; seeger2004gaussian; blei2017variational). Since computing the PPD is often intractable (mackay1992practical), PFNs instead learn to approximate the PPD offline through a process called synthetic prior fitting muller2022transformers. This is achieved using a prior specified by a sampling scheme that first samples a data-generating mechanism, $\phi\sim p(\phi)$, and then samples a synthetic dataset, $D\sim p(D|\phi)$. This process is repeated to generate millions of diverse datasets for training. The network’s parameters, $\theta$, are then optimized to predict held-out test samples ($D_{\text{test}}\subset D$) conditioned on the rest of the dataset ($D_{\text{train}}=D\setminus D_{\text{test}}$). The training objective is to minimize the negative log likelihood (NLL) loss on these held-out examples:
@@
TabPFN is a PFN built specifically for tabular supervised learning (hollmann2023tabpfn). The model employs a novel two-way attention mechanism specifically designed for tabular data. Unlike standard transformers (vaswani2023attentionneed) that treat tabular data as sequential tokens, TabPFN assigns separate representations to each cell in the table. The architecture uses alternating attention patterns: each cell first attends to other features within its row (inter-feature attention), then attends to the same feature across all rows (inter-sample attention). This design ensures permutation invariance for both samples and features while enabling efficient scaling to larger tables than those seen during training. TabPFN v2 (hollmann2025accurate) retains the core training paradigm of the original TabPFN while introducing several key enhancements that improve accuracy, runtime, and applicability. Going forward, when referring to TabPFN, we mean TabPFN v2 since it is the most up-to-date version.
@@ 3 Training TabImpute on Synthetic Data
We develop a featurization for tabular missing data imputation that enables better utilization of TabPFN’s architecture, create a synthetic data generation pipeline across diverse missingness patterns, and employ an adaptive training algorithm to balance performance across all patterns. Training used $8$ H200 GPUs and an Intel Xeon Platinum 8592+ CPU over approximately 2 days, processing $1.9$ million synthetic tables. Our model matches TabPFN’s size and runs on CPU-only systems. Evaluation used $1$ H200 GPU.
@@ 3.1 Entry-wise Featurization and Architecture
We use TabPFN’s base architecture with one modification: removing the attention mask to allow input rows to attend to query rows. Since our input/query row sizes vary randomly with missingness patterns (unlike TabPFN’s controlled synthetic generation), we remove the mask to enable parallel batch training. TabPFN’s mask prevents train points from seeing test feature distributions, which is important in general supervised learning. However, our test set is created using data already available to the observed points, thus alleviating any data-leakage concerns.